please solve the simultaneous equations
5x-2y=13
7x+4y=8

Answer by FrankM(1040)

5x-2y=13
7x+4y=8
The fastest way for me is to simply double both sides of first equation;
10x-4y=26 then add the second equation to this .
17x+0y=34
x=2, so 5(2)-2y=13, 10-2y=13, -2y=3, and last, y = -1.5
check by substituting into second equation -
7(2)+4(-1.5)=8
14 - 6 = 8 (yes!)